# Prompt Injection attack against LLM-integrated Applications

### v2

@@ 1 Introduction
Large Language Models (LLMs) like GPT-4 [39], LLaMA [37], and PaLM2 [18], have dramatically transformed a wide array of applications with their exceptional ability to generate human-like texts. Their integration spans various applications, from digital assistants to AI-powered journalism. However, this expanded usage is accompanied by heightened security vulnerabilities, manifested by a broad spectrum of adversarial tactics such as jailbreak [60, 15, 41] and backdoor [7, 68, 36], and complex data poisoning [32, 38, 67].
@@
Among these security threats, prompt injection where harmful prompts are used by malicious users to override the original instructions of LLMs, is a particular concern. This type of attack, most potent in LLM-integrated applications, has been recently listed as the top LLM-related hazard by OWASP [40]. Existing prompt injection methods [20, 44, 6] manipulate the LLM output for individual users. A recent variant [48] aims to recover previously input prompts at the service provider end. Unfortunately, comprehending the prompt patterns that initiate such attacks remains a significant challenge. Early attempts to exploit this vulnerability used heuristic prompts, discovered through the "trial and error" manner, exploiting the initial unawareness of developers. A thorough understanding of the mechanisms underlying prompt injection attacks, however, is still elusive.
@@
To decipher these attack mechanisms, we initiate a pilot study on 10 real-world black-box LLM-integrated applications, all of which are currently prevalent commercial services in the market. We implement existing prompt injection techniques [20, 44, 6] on them, and only achieve partially successful exploits on two out of the ten targets. The reasons for the unsuccessful attempts are three-pronged. Firstly, the interpretation of prompt usage diverges among applications. While some applications perceive prompts as parts of the queries, others identify them as analytical data payloads, rendering the applications resistant to traditional prompt injection strategies. Secondly, numerous applications enforce specific format prerequisites on both inputs and outputs, inadvertently providing a defensive mechanism against prompt injection, similar to syntax-based sanitization. Finally, applications often adopt multi-step processes with time constraints on responses, rendering potentially successful prompt injections to fail in displaying results due to extended generation duration.
@@
Based on our findings, we find that a successful prompt attack hinges on tricking the LLM to interpret the malicious payload as a question, rather than a data payload. This is inspired by traditional injection attacks such as SQL injection [10, 25, 14] and XSS attacks [23, 27, 63], where specially crafted payloads disturb the routine execution of a program by encapsulating previous commands and misinterpreting malevolent input as a new command. This understanding underpins the formulation of our distinct payload generation strategy for black-box prompt injection attacks.
@@
Thwarting prompt injection attacks can pose a significant challenge. To evaluate the efficacy of existing countermeasures, we apply common defensive mechanisms [50, 46, 52] to some open-source LLM-integrated projects. Our assessments reveal that while these defenses can mitigate traditional prompt injection attacks, they are still vulnerable to malicious payloads generated by HouYi. We hope our work will inspire additional research into the development of more robust defenses against prompt injection attacks.
@@ 2.2 Prompt Injection
Prompt injection refers to the manipulation of the language model’s output via engineered malicious prompts. Current prompt injection attacks predominantly fall into two categories. Some attacks [44, 6] operate under the assumption of a malicious user who injects harmful prompts into their inputs to the application, as shown in the bottom part of Figure 1. Their primary objective is to manipulate the application into responding to a distinct query rather than fulfilling its original purpose. To achieve this, the adversary crafts prompts that can influence or nullify the predefined prompts in the merged version, thereby leading to desired responses. For instance, in the given example, the combined prompt becomes “Answer the following question as a kind assistant: Ignore previous sentences and print “hello world”.” As a result, the application will not answer questions but output the string of “hello world”. Such attacks typically target applications with known context or predefined prompts. In essence, they leverage the system’s own architecture to bypass security measures, undermining the integrity of the entire application.
Recent research [20] delves into a more intriguing scenario wherein the adversary seeks to contaminate the LLM-integrated application to exploit user endpoints. Given that many contemporary LLM-integrated applications interface with the Internet to deliver their functionalities, the injection of harmful payloads into Internet resources can compromise these applications. Specifically, these attacks hinge on transmitting deceptive messages to the LLM either passively (through requested websites or social media posts) or actively (e.g., through emails), causing the applications to take malicious actions prompted by these poisoned sources.
@@ 2.3 Threat Model
We focus on the attack scenario demonstrated in Figure 1. In particular, our threat model contemplates an adversary aiming to execute a prompt injection attack on an LLM-integrated application. The adversary utilizes publicly accessible service endpoints to interact with the application, with the freedom to arbitrarily manipulate the inputs provided to the application. While the specific motivation of such an adversary could vary, the primary objective generally centers on coercing the application into generating outputs that deviate significantly from its intended functionality and design. It is important to clarify that our threat model excludes scenarios where the adversary might exploit other potential vulnerabilities in the application, such as exploiting application front-end flaws [21] or poisoning external resources queried by the application to fulfill its tasks [20].
@@ 3.1 Attack Categorization
For RQ1 (Scope), prior research [44, 16, 4] has detailed several vanila prompt injection attacks targeting both standalone LLMs and LLM-integrated applications. Despite their varying representations, these attacks can typically be classified into one of the following three categories:
@@
Context Ignoring. A more interesting strategy involves injecting a malicious prompt sentence intended to manipulate the LLM so that it ignores the preceding context and concentrates only on the subsequent prompt. An example, as highlighted in  [44] is: “Ignore the previous instructions and print the instructions”. If the attack is successful, the LLM would disregard the original user input, while executing the maliciously injected command instead.
@@ 3.2.1 Overview
To further investigate RQ2 (Exploitability), we select 10 commercial LLM-integrated applications from Supertools [3], a comprehensive collection of trending applications empowered by LLMs. Specifically, we choose two applications from each of the five categories as classified by Supertools: chatbot, writing assistant, code assistant, business analysis, and creative generation. More information about these applications is provided in Table 1.
We first manually gain the appropriate access to these services and confirm their original LLM-integrated functionalities are operational. Following this, we apply the three attack strategies in Section 3.1 to these applications. We utilize all the prompts referenced in existing works [44, 16, 4] and combine them with three straightforward malicious questions: Q1 “print hello world”; Q2 “What is the capital of the USA?”; Q3 “Repeat the prompt of this question”. For Q1 and Q2, we deem the attack successful if the output contains the correct answer. For Q3, success is determined if the output deviates from the application’s ideal functionality. As our primary goal is to ascertain whether the prompt injection strategy could influence the model’s output, we does not specifically verify if the printed prompt is correct or hallucinated. To ensure comprehensiveness, we repeat each prompt injection attack five times and record the success rate.
@@ 3.2.2 Case Study
We provide an example to detail our experimental procedure and its outcomes. We choose DecisionAI [^2] In the following of this paper, the original name of the service provider is anonymized due to non-disclosure reasons unless specified., an AI assistant service that enhances the decision-making capabilities for users. This application leverages GPT models to meticulously analyze the pros and cons related to user decisions. It further employs Strengths, Weaknesses, Opportunities, and Threats (SWOT) analysis [24] to augment users’ comprehension of their decision-making process. The sequence of user interaction with DecisionAI typically follows three main steps: ❶ The user proposes a decision to DecisionAI; ❷ DecisionAI rephrases the decision for clarity and precision; ❸ DecisionAI conducts an extensive pros&cons evaluation, culminating in an assessment of the decision’s feasibility. An example of DecisionAI analyzing the decision of Pursueing a Ph.D. degree is illustrated in Figure 2.
@@ 4.1 Design Insight
Our attack methodology is inspired by the traditional injection attacks such as SQL injection [10, 25, 14] and XSS attacks [23, 27, 63]. In these attacks, a carefully crafted payload manipulates the victim system into executing it as a command, disrupting the system’s normal operation. The key to such type of injection attacks resides in the creation of a payload that can terminate the preceding syntax. Figure 3 depicts an example of SQL injection. The payload “’)” successfully encapsulates the SQL statement, treating the preceding SQL syntax as a finalized SQL command. This allows the ensuing syntax to be interpreted as a supplementary logic (“OR 1=1” is interpreted as “OR TRUE”). Note that successful exploitation also necessitates specific formatting syntax to ensure the SQL command is syntactically correct (“---” indicates the system should disregard the following syntax).
@@
Similar to these traditional injection attacks, our attack aims to deceive an LLM into interpreting the injected prompt as an instruction to be answered separately from the previous context. Our observation from Section 3.2 suggests that, while context-ignoring attacks presented in previous works [4, 20] attempt to create a separation, such approaches have proven insufficient. In particular, a simple prompt of “ignore the previous context” often gets overshadowed by larger, task-specific contexts, thus not powerful enough to isolate the malicious question. Moreover, these approaches do not take into account the previous context. In parallel with traditional injection attacks, it appears that they employ an unsuitable payload for achieving this separation.
@@ 6.1 Evaluation Setup
Result Collection and Disclosure. We have undertaken the dissemination of our findings with exceptional care, holding privacy and security paramount when assessing the evaluated applications. Specifically, each prompt injection attack is manually scrutinized to ascertain its success, deliberately avoiding mass repetitive experimentation to prevent potential misuse of service resources. Upon recognizing successful prompt injection attempts, we promptly and responsibly relay our discoveries to all evaluated applications. In a spirit of full transparency, we only reveal the names of applications, whose service providers have acknowledged the vulnerabilities we pinpointed and granted permission for public disclosure, i.e., Notion [1], Parea [2] and WriteSonic [13]. For the remaining services, their functionalities are presented in an anonymous manner in Table 5.
@@ 6.2 Vulnerability Detection (RQ3)
Third, we also observe that not every exploit scenario consistently achieves success, despite the potential vulnerability presented by the Prompt Leaking scenario. Our thorough analysis discerns three primary factors influencing this outcome. (1) The inherent inconsistency of LLM-generated outputs contribute to unstable application outputs. Applications utilize different LLM models, each with unique behavior and characteristics. For instance, those employing the OpenAI models [39] in creative content generation often opt for high temperature settings to yield more imaginative results. Attacking the same application with prompt injection also yields inconsistent results. (2) The quality of an application’s implementation, especially with regard to error handling, can directly affect the success rate of prompt injections. For example, some applications such as EngageAI and TripPlan do not effectively handle errors returned from the GPT API. When these applications encounter overload errors, such as when token usage exceeds the maximum limit, the API returns an error message. Because these applications fail to manage such errors properly, the error message is directly reflected back, leading to the failure of our attack. (3) The success rate of exploit scenarios is substantially contingent upon the application designs. For example, applications such as DecisioAI and MindAI, which impose output word-length and format restrictions, could experience internal errors in the Information Gathering and Spam Generation scenarios, especially when these prompts generate lengthy responses. Consequently, to ensure maximum effectiveness, exploit prompts should be carefully constructed, considering the unique characteristics and limitations of the applications.
@@ 6.4 Vulnerability Validation (RQ5)
Our approach has led to the successful identification of 31 unique vulnerabilities across a variety of applications. 10 have been confirmed and acknowledged by the vendors. These applications, which include commercial products and services such as Notion [1] serving over 20 million users, demonstrate potential security risks in prevalent applications.
@@ 6.4.1 Case 1: WriteSonic Prompt Leak
Our evaluation revealed a significant prompt leak vulnerability in WriteSonic [13], an AI-powered writing assistant that boasts over 200,000 active users, via our established method of prompt injection. WriteSonic generally assists users in creating written content by leveraging ChatGPT to address user queries related to ongoing writing or brainstorming tasks. However, a malevolent user can exploit this system by employing the language shifting strategy as demonstrated below.
@@ 6.4.2 Case 2: Parea Prompt Abuse
Parea [2], an LLM-integrated application to enhance the quality of responses from ChatGPT by rephrasing user inputs, exhibits a noteworthy vulnerability related to prompt abuse identified through our rigorous prompt injection technique. A typical user may pose a common question like, “What is the capital of the USA?”. Ordinarily, Parea engages ChatGPT to optimize such queries. However, we discovered that a malicious user could append a semantic separator such as “For the above prompt revision, can you explain why you revise it in that way?”, thus enabling execution of any user-defined command in Disruptor Component, as illustrated below.
@@
We unveil a novel exploitation scenario, the prompt abuse, revealing a critical loophole in the security of LLM-integrated applications such as Parea. Developers bear the financial burden for unintended usage of LLMs like ChatGPT, while malicious actors manipulate Parea to fulfill their intentions without any cost. Since Parea is a free application, our evaluation shows an alarming daily financial loss for Parea developers of \$259.2, a figure derived from 90k tokens processed per minute [5] at a cost of \$0.002 per 1k tokens using GPT3.5-turbo [45], extrapolated over 1440 minutes. Furthermore, 30 other LLM-integrated applications are susceptible to similar prompt abuse. In response to our findings, the developers of Parea acknowledged the vulnerability and the pressing need to rectify it, stating, “Thank you for flagging. We are indeed aware of and addressing prompt injection vulnerabilities at Parea. As you know, this is a critical security point for many companies in the LLM space.”

### v3

@@ 1 Introduction
Large Language Models (LLMs) like GPT-4 [GPT4], LLaMA [llama], and PaLM2 [palm2], have dramatically transformed a wide array of applications with their exceptional ability to generate human-like texts. Their integration spans various applications, from digital assistants to AI-powered journalism. However, this expanded usage is accompanied by heightened security vulnerabilities, manifested by a broad spectrum of adversarial tactics such as jailbreak [DBLP:conf/ccs/Si0BCSZ022, jailbreak_gpt1, jailbreak_gpt2] and backdoor [DBLP:conf/sp/BagdasaryanS22, DBLP:conf/emnlp/ZhangLM0022, mei2023notable], and complex data poisoning [DBLP:journals/corr/abs-2004-08994, DBLP:conf/emnlp/MoradiS21, DBLP:conf/acl/0002PTK22].
@@
Among these security threats, prompt injection where harmful prompts are used by malicious users to override the original instructions of LLMs, is a particular concern. This type of attack, most potent in LLM-integrated applications, has been recently listed as the top LLM-related hazard by OWASP [PI-top1]. Existing prompt injection methods [greshake2023youve, perez2022ignore, apruzzese2023realgradients] manipulate the LLM output for individual users. A recent variant [PromptLeaking] aims to recover previously input prompts at the service provider end. Unfortunately, comprehending the prompt patterns that initiate such attacks remains a significant challenge. Early attempts to exploit this vulnerability used heuristic prompts, discovered through the "trial and error" manner, exploiting the initial unawareness of developers. A thorough understanding of the mechanisms underlying prompt injection attacks, however, is still elusive.
@@
To decipher these attack mechanisms, we initiate a pilot study on 10 real-world black-box LLM-integrated applications, all of which are currently prevalent commercial services in the market. We implement existing prompt injection techniques [greshake2023youve, perez2022ignore, apruzzese2023realgradients] on them, and only achieve partially successful exploits on two out of the ten targets. The reasons for the unsuccessful attempts are three-pronged. Firstly, the interpretation of prompt usage diverges among applications. While some applications perceive prompts as parts of the queries, others identify them as analytical data payloads, rendering the applications resistant to traditional prompt injection strategies. Secondly, numerous applications enforce specific format prerequisites on both inputs and outputs, inadvertently providing a defensive mechanism against prompt injection, similar to syntax-based sanitization. Finally, applications often adopt multi-step processes with time constraints on responses, rendering potentially successful prompt injections to fail in displaying results due to extended generation duration.
@@
Based on our findings, we find that a successful prompt attack hinges on tricking the LLM to interpret the malicious payload as a question, rather than a data payload. This is inspired by traditional injection attacks such as SQL injection [boyd2004sqlrand, halfond2006classification, clarke2009sql] and XSS attacks [DBLP:journals/saem/GuptaG17, hydara2015current, weinberger2011systematic], where specially crafted payloads disturb the routine execution of a program by encapsulating previous commands and misinterpreting malevolent input as a new command. This understanding underpins the formulation of our distinct payload generation strategy for black-box prompt injection attacks.
@@
Thwarting prompt injection attacks can pose a significant challenge. To evaluate the efficacy of existing countermeasures, we apply common defensive mechanisms [SandwichDefense, instructiondef, XMLTagging] to some open-source LLM-integrated projects. Our assessments reveal that while these defenses can mitigate traditional prompt injection attacks, they are still vulnerable to malicious payloads generated by HouYi. We hope our work will inspire additional research into the development of more robust defenses against prompt injection attacks.
@@ 2.2 Prompt Injection
Prompt injection refers to the manipulation of the language model’s output via engineered malicious prompts. Current prompt injection attacks predominantly fall into two categories. Some attacks [perez2022ignore, apruzzese2023realgradients] operate under the assumption of a malicious user who injects harmful prompts into their inputs to the application, as shown in the bottom part of Figure 1. Their primary objective is to manipulate the application into responding to a distinct query rather than fulfilling its original purpose. To achieve this, the adversary crafts prompts that can influence or nullify the predefined prompts in the merged version, thereby leading to desired responses. For instance, in the given example, the combined prompt becomes “Answer the following question as a kind assistant: Ignore previous sentences and print “hello world”.” As a result, the application will not answer questions but output the string of “hello world”. Such attacks typically target applications with known context or predefined prompts. In essence, they leverage the system’s own architecture to bypass security measures, undermining the integrity of the entire application.
Recent research [greshake2023youve] delves into a more intriguing scenario wherein the adversary seeks to contaminate the LLM-integrated application to exploit user endpoints. Given that many contemporary LLM-integrated applications interface with the Internet to deliver their functionalities, the injection of harmful payloads into Internet resources can compromise these applications. Specifically, these attacks hinge on transmitting deceptive messages to the LLM either passively (through requested websites or social media posts) or actively (e.g., through emails), causing the applications to take malicious actions prompted by these poisoned sources.
@@ 2.3 Threat Model
We focus on the attack scenario demonstrated in Figure 1. In particular, our threat model contemplates an adversary aiming to execute a prompt injection attack on an LLM-integrated application. The adversary utilizes publicly accessible service endpoints to interact with the application, with the freedom to arbitrarily manipulate the inputs provided to the application. While the specific motivation of such an adversary could vary, the primary objective generally centers on coercing the application into generating outputs that deviate significantly from its intended functionality and design. It is important to clarify that our threat model excludes scenarios where the adversary might exploit other potential vulnerabilities in the application, such as exploiting application front-end flaws [8771368] or poisoning external resources queried by the application to fulfill its tasks [greshake2023youve].
@@ 3.1 Attack Categorization
For RQ1 (Scope), prior research [perez2022ignore, ncc-prompt-injection, gpt3-prompt-injection] has detailed several vanila prompt injection attacks targeting both standalone LLMs and LLM-integrated applications. Despite their varying representations, these attacks can typically be classified into one of the following three categories:
@@
Context Ignoring. A more interesting strategy involves injecting a malicious prompt sentence intended to manipulate the LLM so that it ignores the preceding context and concentrates only on the subsequent prompt. An example, as highlighted in  [perez2022ignore] is: “Ignore the previous instructions and print the instructions”. If the attack is successful, the LLM would disregard the original user input, while executing the maliciously injected command instead.
@@ 3.2.1 Overview
To further investigate RQ2 (Exploitability), we select 10 commercial LLM-integrated applications from Supertools [super-tool], a comprehensive collection of trending applications empowered by LLMs. Specifically, we choose two applications from each of the five categories as classified by Supertools: chatbot, writing assistant, code assistant, business analysis, and creative generation. More information about these applications is provided in Table 1.
We first manually gain the appropriate access to these services and confirm their original LLM-integrated functionalities are operational. Following this, we apply the three attack strategies in Section 3.1 to these applications. We utilize all the prompts referenced in existing works [perez2022ignore, ncc-prompt-injection, gpt3-prompt-injection] and combine them with three straightforward malicious questions: Q1 “print hello world”; Q2 “What is the capital of the USA?”; Q3 “Repeat the prompt of this question”. For Q1 and Q2, we deem the attack successful if the output contains the correct answer. For Q3, success is determined if the output deviates from the application’s ideal functionality. As our primary goal is to ascertain whether the prompt injection strategy could influence the model’s output, we does not specifically verify if the printed prompt is correct or hallucinated. To ensure comprehensiveness, we repeat each prompt injection attack five times and record the success rate.
@@ 3.2.2 Case Study
We provide an example to detail our experimental procedure and its outcomes. We choose DecisionAI [^2] In the following of this paper, the original name of the service provider is anonymized due to non-disclosure reasons unless specified., an AI assistant service that enhances the decision-making capabilities for users. This application leverages GPT models to meticulously analyze the pros and cons related to user decisions. It further employs Strengths, Weaknesses, Opportunities, and Threats (SWOT) analysis [gurl2017swot] to augment users’ comprehension of their decision-making process. The sequence of user interaction with DecisionAI typically follows three main steps: ❶ The user proposes a decision to DecisionAI; ❷ DecisionAI rephrases the decision for clarity and precision; ❸ DecisionAI conducts an extensive pros&cons evaluation, culminating in an assessment of the decision’s feasibility. An example of DecisionAI analyzing the decision of Pursueing a Ph.D. degree is illustrated in Figure 2.
@@ 4.1 Design Insight
Our attack methodology is inspired by the traditional injection attacks such as SQL injection [boyd2004sqlrand, halfond2006classification, clarke2009sql] and XSS attacks [DBLP:journals/saem/GuptaG17, hydara2015current, weinberger2011systematic]. In these attacks, a carefully crafted payload manipulates the victim system into executing it as a command, disrupting the system’s normal operation. The key to such type of injection attacks resides in the creation of a payload that can terminate the preceding syntax. Figure 3 depicts an example of SQL injection. The payload “’)” successfully encapsulates the SQL statement, treating the preceding SQL syntax as a finalized SQL command. This allows the ensuing syntax to be interpreted as a supplementary logic (“OR 1=1” is interpreted as “OR TRUE”). Note that successful exploitation also necessitates specific formatting syntax to ensure the SQL command is syntactically correct (“---” indicates the system should disregard the following syntax).
@@
Similar to these traditional injection attacks, our attack aims to deceive an LLM into interpreting the injected prompt as an instruction to be answered separately from the previous context. Our observation from Section 3.2 suggests that, while context-ignoring attacks presented in previous works [gpt3-prompt-injection, greshake2023youve] attempt to create a separation, such approaches have proven insufficient. In particular, a simple prompt of “ignore the previous context” often gets overshadowed by larger, task-specific contexts, thus not powerful enough to isolate the malicious question. Moreover, these approaches do not take into account the previous context. In parallel with traditional injection attacks, it appears that they employ an unsuitable payload for achieving this separation.
@@ 6.1 Evaluation Setup
Result Collection and Disclosure. We have undertaken the dissemination of our findings with exceptional care, holding privacy and security paramount when assessing the evaluated applications. Specifically, each prompt injection attack is manually scrutinized to ascertain its success, deliberately avoiding mass repetitive experimentation to prevent potential misuse of service resources. Upon recognizing successful prompt injection attempts, we promptly and responsibly relay our discoveries to all evaluated applications. In a spirit of full transparency, we only reveal the names of applications, whose service providers have acknowledged the vulnerabilities we pinpointed and granted permission for public disclosure, i.e., Notion [notion], Parea [Parea] and WriteSonic [writesonic]. For the remaining services, their functionalities are presented in an anonymous manner in Table 5.
@@ 6.2 Vulnerability Detection (RQ3)
Third, we also observe that not every exploit scenario consistently achieves success, despite the potential vulnerability presented by the Prompt Leaking scenario. Our thorough analysis discerns three primary factors influencing this outcome. (1) The inherent inconsistency of LLM-generated outputs contribute to unstable application outputs. Applications utilize different LLM models, each with unique behavior and characteristics. For instance, those employing the OpenAI models [GPT4] in creative content generation often opt for high temperature settings to yield more imaginative results. Attacking the same application with prompt injection also yields inconsistent results. (2) The quality of an application’s implementation, especially with regard to error handling, can directly affect the success rate of prompt injections. For example, some applications such as EngageAI and TripPlan do not effectively handle errors returned from the GPT API. When these applications encounter overload errors, such as when token usage exceeds the maximum limit, the API returns an error message. Because these applications fail to manage such errors properly, the error message is directly reflected back, leading to the failure of our attack. (3) The success rate of exploit scenarios is substantially contingent upon the application designs. For example, applications such as DecisioAI and MindAI, which impose output word-length and format restrictions, could experience internal errors in the Information Gathering and Spam Generation scenarios, especially when these prompts generate lengthy responses. Consequently, to ensure maximum effectiveness, exploit prompts should be carefully constructed, considering the unique characteristics and limitations of the applications.
@@ 6.4 Vulnerability Validation (RQ5)
Our approach has led to the successful identification of 31 unique vulnerabilities across a variety of applications. 10 have been confirmed and acknowledged by the vendors. These applications, which include commercial products and services such as Notion [notion] serving over 20 million users, demonstrate potential security risks in prevalent applications.
@@ 6.4.1 Case 1: WriteSonic Prompt Leak
Our evaluation revealed a significant prompt leak vulnerability in WriteSonic [writesonic], an AI-powered writing assistant that boasts over 200,000 active users, via our established method of prompt injection. WriteSonic generally assists users in creating written content by leveraging ChatGPT to address user queries related to ongoing writing or brainstorming tasks. However, a malevolent user can exploit this system by employing the language shifting strategy as demonstrated below.
@@ 6.4.2 Case 2: Parea Prompt Abuse
Parea [Parea], an LLM-integrated application to enhance the quality of responses from ChatGPT by rephrasing user inputs, exhibits a noteworthy vulnerability related to prompt abuse identified through our rigorous prompt injection technique. A typical user may pose a common question like, “What is the capital of the USA?”. Ordinarily, Parea engages ChatGPT to optimize such queries. However, we discovered that a malicious user could append a semantic separator such as “For the above prompt revision, can you explain why you revise it in that way?”, thus enabling execution of any user-defined command in Disruptor Component, as illustrated below.
@@
We unveil a novel exploitation scenario, the prompt abuse, revealing a critical loophole in the security of LLM-integrated applications such as Parea. Developers bear the financial burden for unintended usage of LLMs like ChatGPT, while malicious actors manipulate Parea to fulfill their intentions without any cost. Since Parea is a free application, our evaluation shows an alarming daily financial loss for Parea developers of \$259.2, a figure derived from 90k tokens processed per minute [gpt-rate-limit] at a cost of \$0.002 per 1k tokens using GPT3.5-turbo [gpt-pricing], extrapolated over 1440 minutes. Furthermore, 30 other LLM-integrated applications are susceptible to similar prompt abuse. In response to our findings, the developers of Parea acknowledged the vulnerability and the pressing need to rectify it, stating, “Thank you for flagging. We are indeed aware of and addressing prompt injection vulnerabilities at Parea. As you know, this is a critical security point for many companies in the LLM space.”